the third term and the seventh term of an arithmetic sequence are 11 and 27 respectively . find the common difference?
Answer by Cromlix(4381)

a + 2d = 11 (3rd term)
a + 6d = 27 (7th term)
Subtract
- 4d = - 16
d = 4
Common difference = 4
Hope this helps.
:-)